# A Novel Hierarchical Co-Optimization Framework for Coordinated Task Scheduling and Power Dispatch in Computing Power Networks

### v2

@@ Abstract
The proliferation of large-scale artificial intelligence and data-intensive applications has spurred the development of Computing Power Networks (CPNs), which deliver ubiquitous and on-demand computational resources. However, the immense energy consumption poses a significant sustainability challenge. Simultaneously, power grids are grappling with the instability by the high penetration of intermittent renewable energy sources (RES). This paper addresses these dual challenges through a novel Two-Stage Co-Optimization (TSCO) framework that synergistically manages power system dispatch and CPN task scheduling to achieve low-carbon operations. The framework decomposes the complex, large-scale problem into a day-ahead stochastic unit commitment stage and a real-time operational stage. The former is solved using Benders decomposition for computational tractability, while in the latter, economic dispatch of generation assets is coupled with an adaptive CPN task scheduling managed by a deep reinforcement learning agent. It makes carbon-aware decisions by responding to dynamic grid conditions, including real-time electricity prices and marginal carbon intensity. Through extensive simulations on an IEEE 30-bus system integrated with a CPN, the TSCO framework is shown to significantly outperform baseline approaches. It can reduce carbon emissions and operational costs, while simultaneously decreasing RES curtailment by more than $60\%$ and maintaining a stringent quality of service for computational tasks.
@@ I-A Background
The digital transformation of the global economy is fueling an unprecedented demand for computational power. The rise of artificial intelligence (AI), big data analytics, and the Internet of Things (IoT) has pushed traditional centralized cloud computing architectures to their limits [1]. In response, a new paradigm known as the Computing Power Network (CPN) has emerged [2]. A CPN aims to interconnect vast, geographically distributed, and heterogeneous computing resources, spanning from large-scale data centers to edge nodes, into a unified, programmable fabric [3]. The core vision of CPN is to break down resource silos and enable the flexible, on-demand scheduling and allocation of computing, storage, and network resources, thereby improving efficiency and enabling novel, low-latency applications.
Concurrently, the global energy sector is undergoing a profound transition towards sustainability, marked by massive investments in Renewable Energy Sources (RES) such as wind and solar power [4]. This shift presents a powerful, yet challenging, symbiotic relationship with the burgeoning CPN paradigm. On one hand, CPNs are voracious consumers of electricity; data centers alone are projected to account for a significant portion of global electricity load growth, with AI workloads being a primary driver. Globally, data centers already account for $1-2\%$ of total electricity consumption, a figure comparable to the aviation industry. Driven by the explosive growth of AI, this demand is projected to double by 2030 [5]. In the United States, data center electricity usage is forecast to climb from $4.4\%$ of the national total in 2023 to as high as $12\%$ by 2028111https://www.energy.gov/articles/doe-releases-new-report-evaluating-increase-electricity-demand-data-centers, with AI workloads shifting from a minor component to a primary driver of this expansion. The availability of clean, renewable energy offers a direct path to decarbonize this massive computational infrastructure[6]. On the other hand, the very nature of CPNs, with their inherent flexibility in workload scheduling, presents a unique opportunity to support and stabilize a grid increasingly reliant on intermittent RES [7].
@@ I-B Research Motivation
The central challenge lies in a fundamental misalignment between the operational dynamics of CPNs and RES-dominated power grids. CPNs are designed to host a wide array of computational tasks, from latency-sensitive services like cloud virtual reality to compute-intensive, batch-processing jobs like large model training [8], [9]. These tasks demand a highly reliable and stable power supply to ensure Quality of Service (QoS). However, RES are inherently variable, non-dispatchable, and dependent on weather conditions. This intermittency introduces significant volatility into the power grid, leading to challenges in maintaining frequency and voltage stability, and often results in the wasteful curtailment of clean energy when generation exceeds demand [10].
Simply connecting a CPN to a grid with high RES penetration without intelligent coordination creates a direct conflict. During periods of low RES output, the CPN’s demand would force the grid to rely on expensive and carbon-intensive fossil-fuel peaker plants to maintain balance. Conversely, during periods of high RES output, the grid may be forced to curtail wind or solar generation to prevent overload, even as CPNs continue to draw power from a mixed-carbon source. This decoupled operation leads to a suboptimal outcome: either unreliable computation or high operational costs and carbon emissions. The problem is not merely about sourcing green energy, but about managing the flexibility of CPN workloads as a grid-stabilizing asset. The inherent ability to shift CPN tasks in time (delaying non-critical jobs) and space (migrating workloads to regions with abundant RES) constitutes a powerful form of Demand Response (DR) [11]. This reframes the problem from “how to power the CPN cleanly” to “how to leverage the CPN’s flexibility to enable a cleaner, more stable grid”. This symbiotic relationship is the core motivation for the co-optimization framework proposed in this paper.
@@ I-C Contributions of the Paper
This paper proposes a holistic, hierarchical co-optimization framework that treats the power system and the CPN as a single, integrated entity. By breaking down the silos between power system operators and CPN schedulers, the framework unlocks significant economic and environmental benefits. It coordinates power generation dispatch with computational task scheduling across multiple timescales to align CPN energy demand with the availability of low-cost, low-carbon renewable energy. As far as we know, this is one of the first efforts to jointly optimize the power grid and the CPN. The main contributions of this work are summarized as follows:
  • A Comprehensive Integrated System Model: A detailed model is developed that captures the intricate interplay between a heterogeneous CPN and a modern power grid. It incorporates conventional thermal generators, stochastic RES, and battery energy storage systems (BESS), while explicitly modeling the spatio-temporal dynamics of electricity prices and carbon intensity.
  • A Novel Two-Stage Co-Optimization (TSCO) Framework: A hierarchical framework is designed to decompose the computationally intractable joint optimization problem into two manageable stages. A day-ahead planning stage addresses long-term unit commitment and resource reservation, while a real-time operational stage handles dynamic economic dispatch and adaptive task scheduling.
@@
  • Adaptive Real-Time Scheduling with Deep Reinforcement Learning (DRL): A DRL agent is developed for the real-time CPN task scheduling subproblem. This model-free approach enables fast, adaptive, and carbon-aware scheduling decisions in response to the highly dynamic and complex state of the joint CPN-grid system, complementing the model-based optimization of the power dispatch.
  • Extensive High-Fidelity Performance Evaluation: The efficacy of the proposed TSCO framework is validated through extensive simulations using realistic data. The evaluation demonstrates significant improvements in carbon emissions, operational costs, and RES utilization when compared against decoupled and carbon-agnostic baseline strategies.
@@ I-D Paper Structure
The remainder of this paper is structured as follows. Section II provides a critical review of related work in CPN scheduling, integrated energy systems, and carbon-aware computing. Section III presents the detailed mathematical formulation of the CPN and power system models, defining the joint optimization problem. Section IV describes the proposed TSCO framework, including the Benders decomposition algorithm and the DRL-based scheduler. Section V details the simulation setup and presents a comprehensive performance evaluation against several baseline methods. Finally, Section VI concludes the paper with a summary of findings and directions for future research.
@@ II Related Works
This section provides a structured review of the literature across three key domains that intersect in this work, including Computing Power Network (CPN) architectures, integrated energy system management, and carbon-aware workload scheduling. This analysis serves to contextualize our contribution and highlight the research gap that our proposed framework aims to fill.
@@ II-A Computing Power Network Architectures and Scheduling
A significant body of research in CPNs has concentrated on task scheduling [12]. The primary objectives have traditionally been to optimize QoS metrics. For instance, studies have focused on developing scheduling policies to minimize task completion delay and enhance reliability, often formulating the problem as a Continuous-Time Markov Decision Process (CMDP) and solving it with DRL techniques. Other works have explored task offloading in terminal-side CPNs [13] or the joint selection of routing paths and computing nodes [14]. While some research considers energy consumption as a constraint or a secondary objective, the explicit, primary optimization of carbon emissions based on the real-time state of the power grid remains largely unexplored. Existing CPN scheduling literature typically treats the power grid as an external, static entity, overlooking the potential for dynamic, symbiotic interaction.
@@ II-B Integrated Energy System Management
The integration of flexible loads and DR has been identified as a key enabler for grids with high RES penetration. Studies have explored how to coordinate data centers as flexible loads with a load aggregator to minimize electricity costs and absorb grid volatility [18]. However, these studies often rely on simplified models of the flexible load, such as an abstract ability to shift power consumption in time, without capturing the complex internal constraints, dependencies, and heterogeneous resource requirements of a CPN workload. They treat the CPN as a “closed box” load, missing the opportunity to optimize its internal operations in concert with the grid [19].
@@ II-C Carbon-Aware Workload Scheduling
The green computing field has produced a substantial body of work on carbon-aware scheduling for geographically distributed data centers. A common strategy is spatio-temporal scheduling [20], which involves shifting computational workloads in time or space to data centers with lower electricity prices or cleaner energy mixes. These methods often leverage real-time carbon intensity signals from services like WattTime [^2] https://watttime.org/data-science/data-signals/ to guide scheduling decisions.
To manage carbon emissions over the long term, some works have proposed online algorithms based on Lyapunov optimization [21]. This technique transforms a long-term average constraint, e.g., a carbon budget, into a series of real-time optimization subproblems by maintaining a “virtual queue” that tracks the deviation from the budget. The scheduler is then penalized for actions that increase this queue length [22]. While powerful, this line of research suffers from a critical limitation. It almost universally treats the power grid as an exogenous system [23]. The price and carbon intensity signals are assumed to be external inputs that are unaffected by the scheduling decisions. This assumption breaks down at scale, as the collective actions of large CPNs can and will influence grid operations, market prices, and the generation mix, thereby altering the very signals they are responding to.
@@ II-D Research Gap Summary
  • Simplifies the CPN: Power system and DR studies model flexible loads like data centers in an overly simplistic manner, failing to capture the rich internal complexity of CPN workloads, resource heterogeneity, and QoS constraints.
This leads to a new class of control problem where the decision variables are distributed across two separate domains. The system dynamics are characterized by a mix of well-understood physics (the power grid) and complex, stochastic behavior (the CPN). A model-based optimization approach is ill-suited for the fast, dynamic CPN scheduling, while a model-free AI approach cannot guarantee adherence to the hard physical constraints of the power grid. Consequently, there is a clear need for a hybrid framework that endogenously models the bi-directional interactions between the two systems. Our TSCO framework, which combines large-scale, model-based optimization for slow, physics-heavy planning with a model-free, adaptive AI technique for fast, complex real-time scheduling, is designed specifically to fill this gap.
@@ III System Architecture and Problem Formulation
This section presents the mathematical models for the CPN and the integrated power system. These models form the basis of our co-optimization framework, as shown in Fig. 1. For clarity, Table I summarizes the key notations used throughout this paper.
@@
TABLE I: Summary of Key Notations
Symbol Description
$\mathcal{N}$ Set of CPN nodes (data centers)
$\mathcal{G}_{C}$ Set of conventional thermal generators
$\mathcal{G}_{R}$ Set of renewable energy (RES) generators (wind/solar)
$\mathcal{J}$ Set of computational jobs in CPN
$\mathcal{T}_{k}$ Set of subtasks of job $k$
$\mathcal{I}$ Set of buses in the IEEE 30-bus power system
$C_{n,h}^{comp}$ Computing capacity of hardware $h$ at CPN node $n$ (in FLOPS)
$P_{n,t}^{CPN}$ Total power consumption of CPN node $n$ at time $t$
$P_{g,t}$ Power output of conventional generator $g$ at time $t$ (in MWh)
$P_{g,t,\omega}^{R,avail}$ Available power output of RES generator $g$ at time $t$ under scenario $\omega$
$x_{k,\tau,n,h,t}$ Binary variable: 1 if subtask $\tau$ of job $k$ is assigned to hardware $h$ at node $n$ and time $t$
$u_{g,t}$ Commitment status of conventional generator $g$ at time $t$ (1=on, 0=off)
$SOC_{b,t,\omega}$ State-of-charge of battery energy storage system (BESS) $b$ at time $t$ under scenario $\omega$
$\epsilon_{g}$ Carbon emission factor of generator $g$ (tons of $\text{CO}_{2}$ per MWh)
$E_{budget}$ Long-term carbon emission budget for the planning horizon
$LMP_{i,t}$ Locational marginal price at bus $i$ and time $t$ (in \$/MWh)
$MCI_{i,t}$ Marginal carbon intensity at bus $i$ and time $t$
$\Omega$ Set of uncertainty scenarios
@@ III-A Computing Power Network Model
The CPN is modeled as a directed graph $\mathcal{G}_{CPN}=(\mathcal{N},\mathcal{L})$, where $\mathcal{N}$ is the set of geographically distributed CPN nodes, representing data centers, and $\mathcal{L}$ is the set of communication links connecting them.
@@ III-A2 Task Model
Each subtask $\tau\in T_{k}$ is further defined by: (1) hardware type constraint $h_{\tau}\subseteq H$, the set of hardware types capable of executing $\tau$, e.g., large-model training subtasks require $h_{\tau}=\{GPU,TPU\}$; (2) data volume $d_{\tau}$, in GB, for communication constraint calculation; (3) total computational workload $w_{\tau}$ in FLOPs, and resource requirement $r_{\tau,h}$ is the number of processing units for hardware $h\in h_{\tau}$. The execution time of subtask $\tau$ on hardware $h$ of node $n$ is:
@@ III-B Integrated Power System Model
The power grid is modeled based on the IEEE 30-bus test system [^3] https://icseg.iti.illinois.edu/ieee-30-bus-system/, a standard benchmark for power system studies. The system consists of a set of buses $\mathcal{I}$ connected by transmission lines. Each CPN node $n\in\mathcal{N}$ is co-located with a specific load bus $i\in\mathcal{I}$.

### v3

@@ Abstract
The proliferation of large-scale AI and data-intensive applications has driven the development of Computing Power Networks (CPN). It is a key paradigm for delivering ubiquitous, on-demand computational services with high efficiency. However, CPNs face dual challenges in service computing. Immense energy consumption threatens sustainable operations. And the integration with power grids also features high penetration of intermittent Renewable Energy Sources (RES), complicating task scheduling while ensuring Quality of Service (QoS). To address these issues, this paper proposes a novel Two-Stage Co-Optimization (TSCO) framework. It synergistically coordinates CPN task scheduling and power system dispatch, aiming to optimize service performance while achieving low-carbon operations. The framework decomposes the complex, large-scale problem into a day-ahead stochastic unit commitment stage and a real-time operational stage. The former is solved using Benders decomposition for computational tractability, while in the latter, economic dispatch of generation assets is coupled with an adaptive CPN task scheduling managed by a deep reinforcement learning agent. It makes carbon-aware decisions by responding to dynamic grid conditions, including real-time electricity prices and marginal carbon intensity. Extensive simulations demonstrate that the TSCO outperforms baseline approaches significantly. It reduces carbon emissions by $16.2\%$ and operational costs by $12.7\%$, while decreasing RES curtailment by over $60\%$, maintaining a task success rate of $98.5\%$, and minimizing average task tardiness to $12.3$s. This work advances cross-domain service optimization in CPNs.
@@ I-A Background
The digital transformation of the global economy is fueling an unprecedented demand for computational power. The rise of artificial intelligence (AI) has pushed traditional centralized cloud computing architectures to their limits [1]. In response, a new paradigm known as the Computing Power Network (CPN) has emerged [2]. It aims to interconnect vast, distributed, and heterogeneous computing resources into a unified, programmable fabric [3]. The core vision of CPN is to break down resource silos and enable the flexible, on-demand scheduling and allocation of computing, storage, and network resources, thereby enabling novel, low-latency applications. Unlike grid computing or conventional data centers, CPNs explicitly integrate computing task scheduling with power system dispatch.
Concurrently, the global energy sector is undergoing a profound transition towards sustainability, marked by massive investments in Renewable Energy Sources (RES) and Battery Energy Storage Systems (BESS) [4]. This shift presents a symbiotic relationship with the burgeoning CPN paradigm. On the one hand, CPNs are voracious consumers of electricity; data centers alone are projected to account for a significant portion of global electricity load growth, with AI workloads being a primary driver. Globally, data centers already account for $1-2\%$ of total electricity consumption, a figure comparable to the aviation industry. Driven by the explosive growth of AI, this demand is projected to double by 2030 [5]. In the United States, data center electricity usage is forecast to climb from $4.4\%$ of the national total in 2023 to as high as $12\%$ by 2028111https://www.energy.gov/articles/doe-releases-new-report-evaluating-increase-electricity-demand-data-centers, with AI workloads becoming a primary driver of this expansion. The availability of clean, renewable energy offers a direct path to decarbonize this massive computational infrastructure[6]. On the other hand, the very nature of CPNs, with their inherent flexibility in workload scheduling, presents a unique opportunity to support and stabilize a grid increasingly reliant on intermittent RES [7].
@@ I-B Research Motivation
The core challenge lies in a misalignment between the operational dynamics of CPNs and RES-dominated power grids. CPNs are designed to host a wide array of computational tasks, from latency-sensitive services, to compute-intensive, batch-processing jobs [8]. These tasks demand a highly reliable and stable power supply to ensure Quality of Service (QoS). However, RES are inherently variable and dependent on weather conditions. This intermittency introduces significant volatility into the power grid. It often results in the wasteful curtailment of clean energy when generation exceeds demand [9].
Simply connecting a CPN to a grid with high RES penetration without intelligent coordination creates a direct conflict. During periods of low RES output, the CPN’s demand would force the grid to rely on expensive and carbon-intensive fossil-fuel peaker plants to maintain balance. Conversely, during periods of high RES output, the grid may be forced to curtail wind or solar generation to prevent overload, even as CPNs continue to draw power from a mixed-carbon source. This decoupled operation leads to a suboptimal outcome: either unreliable computation or high operational costs and carbon emissions. The problem is not merely about sourcing green energy, but about managing the flexibility of CPN workloads as a grid-stabilizing asset. The inherent ability to shift CPN tasks in time (delaying non-critical jobs) and space (migrating workloads to regions with abundant RES) constitutes a powerful form of Demand Response (DR) [10], [11]. This reframes the problem from “how to power the CPN cleanly” to “how to leverage the CPN’s flexibility to enable a cleaner, more stable grid”. This symbiotic relationship is the core motivation for the co-optimization framework proposed in this paper.
@@ I-C Contributions of the Paper
In the present work, we propose a holistic, hierarchical co-optimization framework that jointly optimizes power system operations and CPN task scheduling. By breaking down the silos between power system operators and CPN schedulers, the framework unlocks significant economic and environmental benefits. It coordinates power generation dispatch with computational task scheduling across multiple timescales to align CPN energy demand with the availability of low-cost, low-carbon renewable energy. To the best of our knowledge, this is one of the first efforts to jointly optimize the power grid and the CPN. The main contributions of this work are summarized next:
  • A Comprehensive Integrated System Model: A detailed model is developed that captures the intricate interplay between a heterogeneous CPN and a modern power grid. It incorporates conventional thermal generators, stochastic RES, and BESS, while explicitly modeling the spatio-temporal dynamics of electricity prices and carbon intensity.
  • A Two-Stage Co-Optimization (TSCO) Framework: A hierarchical framework is designed to decompose the computationally intractable joint optimization problem into two manageable stages. A day-ahead planning stage addresses long-term unit commitment and resource reservation, while a real-time operational stage handles dynamic economic dispatch and adaptive task scheduling.
@@
  • Adaptive Real-Time Scheduling with Deep Reinforcement Learning (DRL): A DRL agent is developed for the real-time CPN task scheduling. This model-free approach enables fast, adaptive, and carbon-aware scheduling decisions in response to the highly dynamic and complex state of the joint CPN-grid system.
@@ I-D Paper Structure
The remainder of the paper is structured as follows. Section II provides a review of related work in CPN scheduling, integrated energy systems, and carbon-aware computing. Section III presents the detailed mathematical formulation of the CPN and power system models, defining the joint optimization problem. Section IV describes the proposed TSCO framework, including the Benders decomposition algorithm and the DRL-based scheduler. Section V details the simulation setup and presents a comprehensive performance evaluation against several baseline methods. Finally, Section VI concludes the paper with a summary of findings and directions for future research.
@@ II-A Computing Power Network Architectures and Scheduling
A significant body of research in CPNs has concentrated on task scheduling [12]. The primary objectives have traditionally been to optimize QoS metrics. For instance, studies have focused on developing scheduling policies to minimize task completion delay and enhance reliability, often formulating the problem as a Continuous-Time Markov Decision Process (CMDP) and solving it with DRL techniques [13]. Other works have explored task offloading in terminal-side CPNs or the joint selection of routing paths and computing nodes [14].
Two other works also considered the collaboration between the CPN and the power grid. Wen et al. [15], [16] proposed a spatiotemporal task scheduling scheme for CPNs, leveraging a customized DRL algorithm to optimize energy consumption and carbon emissions within CPNs. However, their work focuses solely on intra-domain optimization, treating energy supply as an exogenous input without involving power system dynamics. Zhong et al. [17] integrated Electric Vehicles (EVs) into CPN and Power Distribution Network (PDN) coordination. They use EVs as dual carriers of energy storage and computing resources for task offloading. Nevertheless, their CPN scheduling is subordinate to cost minimization of EV-PDN synergy, lacking deep coupling with bulk power systems and carbon-aware optimization. In contrast, our TSCO framework emphasizes bidirectional coupling between CPN and bulk power systems, treating CPN as a flexible load to enhance grid stability and low-carbon operations. Table I compares the approaches with ours.
While some research considers energy consumption as a constraint or a secondary objective, the explicit, primary optimization of carbon emissions based on the real-time state of the power grid remains largely unexplored.
TABLE I: Comparison of CPN and Power Grid Coordination
Key Dimension TSCO (Our Work) [15], [16] [17]
Research Paradigm Cross-domain co-optimization Intra-CPN fine-grained scheduling Multi-agent synergy
Collaborative Entities CPN, bulk power system CPN CPN, PDN, EVs
Optimization Core Low carbon, RES curtailment reduction, grid stability, communication QoS Energy efficiency, task delay, load balancing System cost minimization
System Interaction Bidirectional (CPN and grid) No interaction Unidirectional (CPN/EV constrained by PDN)
Key Innovation Bidirectional coupling; endogenous carbon intensity Triple-selection (nodes/paths/time); intra-CPN coordination EV dual-resource (energy/computing) exploitation
@@ II-B Integrated Energy System Management
The integration of flexible loads and DR has been identified as a key enabler for grids with high RES penetration. Studies have explored how to coordinate data centers as flexible loads with a load aggregator to minimize electricity costs and absorb grid volatility [21]. However, these studies often rely on simplified models of the flexible load, such as an abstract ability to shift power consumption in time, without capturing the complex internal constraints, dependencies, and heterogeneous resource requirements of a CPN workload. They treat the CPN as a “closed box” load, missing the opportunity to optimize its internal operations in concert with the grid.
@@ II-C Carbon-Aware Workload Scheduling
A substantial body of work on carbon-aware scheduling for geographically distributed data centers. A common strategy is spatio-temporal scheduling [22], which involves shifting computational workloads in time or space to data centers with lower electricity prices or cleaner energy mixes. These methods often leverage real-time carbon intensity to guide scheduling.
To manage carbon emissions over the long term, some works have proposed online algorithms based on Lyapunov optimization [23]. This technique transforms a long-term average constraint into a series of real-time optimization subproblems by maintaining a “virtual queue” that tracks the deviation from the budget. The scheduler is then penalized for actions that increase this queue length [24]. This line of research suffers from a critical limitation. It almost universally treats the power grid as an exogenous system. The price and carbon intensity signals are assumed to be external inputs that are unaffected by the scheduling decisions. This assumption breaks down at scale, as the collective actions of large CPNs can influence grid operations, market prices, and the generation mix. Specifically, we have done a study comparing four countries: Germany, the Netherlands, France, and Brazil, as these have very different energy mixes and therefore the variability of the carbon intensity is very different among them.
@@ II-D Research Gap Summary
  • Simplifies the CPN: Power system and DR studies model flexible loads in an overly simplistic manner, failing to capture the rich internal complexity of CPN workloads, resource heterogeneity, and QoS constraints.
This leads to a new class of control problems where the decision variables are distributed across two separate domains. The system dynamics are characterized by a mix of well-understood physics (the power grid) and complex, stochastic behavior (the CPN). A model-based approach is ill-suited for fast, dynamic CPN scheduling, while a model-free AI approach cannot guarantee adherence to hard physical constraints of the power grid. Consequently, there is a need for a novel framework that endogenously models the bidirectional interactions between them. Our TSCO framework, which combines large-scale, model-based optimization for slow, physics-heavy planning with a model-free, adaptive AI technique for fast, complex real-time scheduling, is designed specifically.
@@ III System Architecture and Problem Formulation
These models form the basis of our co-optimization framework, as shown in Fig. 1.
@@ III-A Computing Power Network Model
The CPN is modeled as a directed graph $\mathcal{G}_{CPN}=(\mathcal{N},\mathcal{L})$, where $\mathcal{N}$ is the set of geographically distributed CPN nodes, and $\mathcal{L}$ is the set of communication links connecting them.
@@ III-A2 Task Model
Each subtask $\tau\in T_{k}$ is further defined by: (1) hardware type constraint $h_{\tau}\subseteq H$, the set of hardware types capable of executing $\tau$, e.g., large-model training subtasks require $h_{\tau}=\{GPU,TPU\}$; (2) data volume $d_{\tau}$ in GB, for communication constraint calculation; (3) total computational workload $w_{\tau}$ in FLOPs, and resource requirement $r_{\tau,h}$ is the number of processing units for hardware $h\in h_{\tau}$. The execution time of subtask $\tau$ on hardware $h$ of node $n$ is:
@@ III-B Integrated Power System Model
The system consists of a set of buses $\mathcal{I}$ connected by transmission lines. Each CPN node $n\in\mathcal{N}$ is co-located with a specific load bus $i\in\mathcal{I}$.